# iA: Imperative Learning-based A Search for Path Planning

### v2

@@ IV Experiments
We conduct experiments to evaluate the performance of iA on path planning problems across diverse environments. To ensure a comprehensive comparison, we select the latest data-driven search-based methods as our benchmarks and widely used datasets to provide experimental instances across diverse environments. Furthermore, to assess the performance of the iA algorithm in robot navigation tasks, we conduct simulation experiments in simulated scenarios with a mobile robot. The ablation study focuses on the impact of IL strategy and network choice on network training.
@@ IV-A Datasets and Benchmarks
To provide training data and diverse test instances for evaluation, we select three publicly available datasets representing distinct environmental configurations. Motion Planning Dataset (MP Dataset) [37], Maze Dataset [38], and Matterport3D Dataset [39]. These datasets are widely used for evaluating pathfinding methods. The details about the selected datasets are shown as follows:
  • Motion Planning Dataset (MP Dataset) [37]: This dataset contains a set of binary maps about eight diverse environments with simple obstacle settings.
  • Maze Dataset [38]: This dataset aims to facilitate maze-solving tasks and contains several types of mazes more sophisticated than maps in the MP dataset.
  • Matterport3D Dataset [39]: This dataset provides several Truncated Signed Distance Function (TSDF) maps of indoor scenarios. To get the obstacle setting of these maps, we set a threshold as 0.2 $m$ to select the obstacles.
We use the instance generation mentioned in [8], randomly sampling the start and goal positions in the given maps and generating the instances for training and testing.
To evaluate the path planning performance of the proposed iA framework, we selected the latest open-source data-driven methods, Neural A  [8], Focal Search with Path Probability Map (FS + PPM) [9] and Weighted A + correction factor (WA + CF) [9], as competitors. Besides, we also select the Weighted A [35] as a classical competitor that biases the search toward the goal area to improve search efficiency. The Neural A, FS + PPM, and WA + CF have been trained by the MP dataset in a supervised fashion, given the optimal solution as labels, and we use the well-trained parameters published on their website. For fair comparisons, all the selected data-driven methods are trained exclusively on the MP dataset. The Maze and Matterport3D datasets are used solely for evaluation and remain unseen during training.
IV-B Metrics
  • Reduction ratio of node explorations (Exp) [8]: To measure the search efficiency of the examined method, we utilize the ratio of the reduced search area to the classical A search as a metric, which is defined as
    $\textit{Exp}=100\times\frac{A^{*}-A}{A^{*}},$ (10)
    where $A$ represents the search area of the examined method and $A^{*}$ is the search area of classical A search.
  • Reduction ratio of time (Rt): To measure the time efficiency of the examined method, we utilize the ratio of saved time to the operation time of standard A search as a metric, which is defined as
    $\textit{Rt}=100\times\frac{t^{*}-t}{t^{*}},$ (11)
    where $t^{*}$ is the operating time of classical A solution, and $t$ is the operating time of the examined method.
    $\textit{AL}=\sqrt{\mathcal{L}_{a}(\mu)}+\mathcal{L}_{l}(\mu),$ (12)
    where $\mathcal{L}_{a}(\mu)$ and $\mathcal{L}_{l}(\mu)$ are the search area and path length, given path $\mu$. For consistent unit, we use the square root of $\mathcal{L}_{a}(\mu)$ to plus $\mathcal{L}_{l}(\mu)$ as the final metric.
IV-C Overall Performance
@@
64 A 0 0 93.5 0 0 210.9 0 0 99.2
WA 63.5 47.0 90.2 21.7 27.4 201.0 48.1 36.8 82.4
Neural A 56.0 24.5 95.1 30.9 40.8 206.7 42.3 28.4 86.5
WA+CF 68.8 48.0 90.2 47.3 45.3 200.6 57.2 42.7 80.8
FS+PPM 88.3 83.6 79.3 -8.9 -105.0 214.9 9.3 -43.8 93.3
iA(Ours) 69.4 52.8 86.7 37.8 44.9 198.7 57.8 44.7 78.6
128 WA 70.1 55.4 200.8 69.4 64.4 232.4 68.6 66.0 168.6
Neural A 67.2 42.1 216.8 65.7 55.6 250.0 60.3 53.6 193.7
iA(Ours) 75.9 62.3 192.7 75.2 64.5 227.2 73.7 68.9 161.0
256 WA 60.9 57.9 425.1 89.9 81.8 370.8 59.5 42.7 364.0
Neural A 57,4 52.5 452.4 75.9 65.3 425.5 59.0 30.6 393.8
iA(Ours) 61.9 59.0 423.1 90.4 87.3 370.1 62.2 45.5 355.4
Comparisons with Benchmarks
@@
To evaluate the path planning performance of iA, we conduct both qualitative and quantitative experiments. Fig. 4 visualizes the planning results of the selected methods. From the results, the proposed iA can greatly reduce the search area and still generate paths similar to optimal paths generated by classical A. Furthermore, Table I summarizes the quantitative experimental results across the MP, Maze, and Matterport3D datasets with three different map sizes $64\times 64$, $128\times 128$, and $256\times 256$. To ensure fairness, we conducted 10 randomized trials for each kind of map in all three datasets, sampling start and goal positions uniformly. Final results are the average scores of these trials with the mentioned metrics. Notably, FS + PPM and WA + CF only support the map with a fixed shape, $64\times 64$, because their network designs include fully connected layers that necessitate a predetermined input shape. Adjustments and re-training procedures are required to meet varied input instance shapes. Thus, comparisons at larger map sizes mainly focus on WA, Neural A, and iA.
Map with $\mathbf{64\times 64}$ size
While the FS + PPM algorithm presents better performance in the MP dataset, the FS + PPM algorithm performs significantly worse in the other two datasets, even searching more areas and spending more time than the classical A. For example, the FS + PPM even increases 105% operation time, 8.9% search area, and gets higher AL values for the Maze dataset, compared with the classical A search. Therefore, it is likely that FS + PPM is overfitting on the MP dataset. If we exclude the FS-PPM from consideration, the proposed iA demonstrates superior performance of search efficiency on this dataset among the rest of the selected methods. Besides, iA also has a satisfactory performance in the remaining two datasets, getting the minimum values in the AL metric.
Figure 5: The trajectories of a mobile robot under given navigation tasks in the simulation environments, Tunnel (the first row) and Indoor (the second row). (a) represents the given instances, start (red circle), and goal (green circle). (b), (c) and (d) represent the trajectories of a mobile robot navigating from start to goal in the simulation environments with Classical A, Neural A, and iA.
Map with larger sizes
In larger maps, iA outperforms all the baselines in all metrics, achieving the lowest scores in AL and the highest scores in both Exp and Rt. Specifically, iA averagely reduces 73.2% search area and saves 64.6% search time, demonstrating its superiority in improving search efficiency. Furthermore, iA attains the lowest AL scores across all experiments, highlighting its capacity to balance search area and path length on maps with different sizes.
@@
In summary, the proposed iA gets the lowest AL scores in 8 out of 9 comparisons on diverse environments with different map sizes. Besides, compared with Neural A, iA achieves an average of 11.0% better trade-off between path length and search area, while saving 15.2% runtime and 9.6% search areas, demonstrating iA can effectively reduce the search area while still finding the suboptimal paths. Notably, its consistent performance across different map sizes validates the method’s strong generalization without overfitting, which is critical for robotic applications.
IV-D Autonomous Exploration Development Environments
To validate the practicality, we further evaluate iA and baseline methods in the widely used Autonomous Exploration Development Environment [40] that includes a mobile robot and several scenarios resembling real-world settings. We use the indoor and the tunnel scenarios, shown as Fig. 5 (a), as our test environments. Specifically, we convert their dense point cloud maps into two-dimensional binary maps, where each grid cell represents a $0.5\mathrm{m}\times 0.5\mathrm{m}$ area in the original map. A grid cell will be regarded as an obstacle if its maximum $z$-value exceeds a given threshold; otherwise, it will be regarded as a free cell. Ultimately, the Tunnel and Indoor scenes have a size of $654\times 505$ and $259\times 199$, respectively. Since these map sizes are different, Neural A is the only applicable data-driven baseline method. As a result, we mainly compare the performance of A, Neural A, and iA. It is worth noting that both the Neural A and iA use the MP dataset for training and never see these scenarios.
In tests, the goal position for the mobile robot is manually given. We use the selected methods to generate paths and utilize the integrated path-following method within the maps to provide control commands for navigation. As shown in Fig. 5, both Neural A and iA can generate obstacle-free paths, similar to the optimal path generated by A, and allow the robot to finish the task successfully. As shown in Table II, iA significantly improves the planning efficiency compared to the baseline methods. Specifically, in the Indoor scenario, iA reduces 43.28% search area and saves 52.04% time; in the Tunnel scenario, it reduces 62.73% search area and saves 62.84% time. Besides, in these two scenarios, iA gets lower AL values, compared with Neural A, presenting a better trade-off between search area and path length.
@@ IV-D Autonomous Exploration Development Environments
A Neural A iA
Indoor Exp (Area) 0 (9826) 3.95% (9438) 43.28% (5573)
@@ IV-E Ablation Study
We next evaluate the effectiveness of the IL strategy under different configurations. Four different configurations are considered by combining two search engines (A and WA) and two network architectures (CNN and UNet). Each configuration is trained both with and without the proposed IL training strategy. For non-IL training methods, we follow the Neural A framework, employing the ground truth optimal paths as supervision. All other training settings, including the learning rate and optimizer, are kept consistent to ensure a fair comparison. To assess the trade-off between path optimality and search efficiency, we track the variation of the AL throughout the validation processes and show it in Fig. LABEL:fig:AUNET-LABEL:fig:WACNN, where solid and dashed lines represent the methods with and without IL-based training, respectively. As illustrated, the methods with IL consistently converge to significantly lower AL values than their non-IL counterparts. On average, our IL-based methods achieve approximately a 9.79% reduction in AL, demonstrating the effectiveness of IL and balancing between path optimality and search efficiency.
@@
To further analyze the individual contributions of search engines (A and WA) and network architectures (CNN and UNet) under IL strategy, we isolate and compare the results (solid lines in Fig. LABEL:fig:ablation_study) of the IL-trained configurations. The results show that models using the A search engine (red and blue curves) outperform those employing WA (green and yellow curves) in terms of lower AL scores. Moreover, the UNet-based models present superior performance compared to CNN-based models when paired with the same search engine. Consequently, based on these empirical observations, we adopt the configuration comprising the A search algorithm and the UNet architecture for our final framework.
V Conclusion

### v4

@@ IV Experiments
We conduct experiments to evaluate the performance of iA* on path planning problems across diverse environments. To ensure a comprehensive comparison, we select the latest data-driven search-based methods as our benchmarks and widely used datasets to provide experimental instances across diverse environments. Furthermore, to assess the performance of the iA* algorithm in robot navigation tasks, we conduct simulation experiments in simulated scenarios with a mobile robot. The ablation study focuses on the impact of IL strategy and network choice on network training.
@@ IV-A Datasets and Benchmarks
To provide training data and diverse test instances for evaluation, we select three publicly available datasets representing distinct environmental configurations. Motion Planning Dataset (MP Dataset) [bhardwaj2017learning], Maze Dataset [maze-dataset], and Matterport3D Dataset [chang2017matterport3d]. These datasets are widely used for evaluating path planning methods. The details about the selected datasets are shown as follows:
  • Motion Planning Dataset (MP Dataset) [bhardwaj2017learning]: This dataset contains a set of binary maps about eight diverse environments with simple obstacle settings.
  • Maze Dataset [maze-dataset]: This dataset aims to facilitate maze-solving tasks and contains several types of mazes more sophisticated than maps in the MP dataset.
  • Matterport3D Dataset [chang2017matterport3d]: This dataset provides several Truncated Signed Distance Function (TSDF) maps of indoor scenarios. To get the obstacle setting of these maps, we set a threshold as 0.2 $m$ to select the obstacles.
We use the instance generation mentioned in [yonetani2021nastar], randomly sampling the start and goal positions in the given maps and generating the instances for training and testing.
To evaluate the path planning performance of the proposed iA* framework, we selected the latest open-source data-driven methods, Neural A*  [yonetani2021nastar], Focal Search with Path Probability Map (FS + PPM) [kirilenko2023transpath], and Weighted A* + correction factor (WA* + CF) [kirilenko2023transpath], as competitors. Besides, we also select the Jump Point Search (JPS) [harabor2011jps] and Weighted A* (W*) [pohl1970heuristic] as classical competitors. The JPS algorithm improves the search efficiency by skipping over large sections of nodes through directional pruning and identifying critical jump points, and Weighted A* biases the search toward the goal area to improve search efficiency. The Neural A*, FS + PPM, and WA* + CF have been trained by the MP dataset in a supervised fashion, given the optimal solution as labels, and we use the well-trained parameters published on their website. For fair comparisons, all the selected data-driven methods are trained exclusively on the MP dataset. The Maze and Matterport3D datasets are used solely for evaluation and remain unseen during training. For the training settings, like learning rate and optimizer, we follow the settings in [yonetani2021nastar, kirilenko2023transpath] and empirically adjust them by extensive training trials.
@@
64 A* 0(0) 0(0) 85.9(17.8) 73.0(3.1) 0(0) 0(0) 142.9(31.9) 131.0(31.3) 0(0) 0(0) 42.4(5.2) 37.2(4.6)
JPS 95.1(5.9) - 77.2(22.3) 73.0(3.1) -103(26.7) - 147.9(107.9) 131.0(31.3) -15.2(19.2) - 42.6(18.7) 37.2(4.6)
WA* 64.8(3.0) 47.7(6.5) 82.6(20.7) 75.0(3.7) 43.9(10.4) 33.2(14.0) 144.6(32.2) 135.7(31.1) 48.1(13.8) 36.1(31.3) 42.9(5.2) 39.1(4.7)
Neural A* 63.5(2.6) 39.6(5.4) 86.8(20.4) 79.7(3.7) 40.6(13.1) 37.5(19.2) 151.0(30.9) 142.7(29.7) 42.3(13.4) 10.0(60.2) 44.4(5.7) 40.8(5.4)
WA*+CF 68.8(2.6) 48.0(7.2) 85.6(22.3) 78.7(4.1) 56.8(11.1) 42.7(16.8) 149.2(30.0) 141.5(29.0) 57.2(16.0) 45.6(22.3) 45.6(6.2) 42.1(6.0)
FS+PPM 89.0(0.7) 88.8(0.9) 78.0(16.0) 73.8(3.1) -7.9(10.9) -15.0(31.0) 149.6(30.4) 139.1(29.6) 9.3(26.1) -17.4(56.3) 42.7(5.3) 38.3(5.1)
iA*(Ours) 69.4(2.9) 50.4(5.7) 82.3(19.0) 75.0(3.4) 45.7(13.5) 33.2(17.8) 143.6(32.2) 134.9(30.9) 57.8(13.4) 44.7(30.2) 41.9(5.5) 38.2(4.9)
128 JPS 98.7(2.0) - 155.0(43.4) 147.4(10.0) 47.7(12.7) - 431.0(265.3) 401.3(258.8) -111.9(27.6) - 115.9(77.1) 95.3(13.0)
WA* 65.2(6.2) 51.4(13.0) 167.5(13.1) 151.9(12.0) 52.2(17.3) 35.8(17.0) 432.3(257.1) 412.1(262.5) 67.3(12.9) 56.5(39.3) 106.3(15.0) 100.9(13.4)
Neural A* 67.2(6.6) 46.0(12.3) 173.0(13.4) 158.0(12.1) 59.4(11.7) 42.7(16.9) 442.3(261.7) 423.5(256.1) 71.1(13.0) 55.4(41.7) 116.9(16.4) 111.6(14.7)
iA*(Ours) 67.6(6.1) 56.0(12.2) 166.4(11.3) 151.4(10.0) 54.0(11.8) 39.5(17.6) 427.1(263.0) 406.3(257.8) 74.4(11.5) 60.0(35.9) 105.9(14.9) 100.6(13.4)
256 JPS 99.4(1.5) - 300.4(83.2) 277.9(25.8) 56.8(8.1) - 744.6(444.2) 703.4(426.7) -204.1(526) - 207.5(162.4) 153.2(28.0)
WA* 74.1(16.3) 62.5(28.3) 340.5(60.0) 300.0(47.5) 63.9(16.7) 31.2(36.0) 741.5(435.6) 714.8(423.4) 70.0(12.0) 54.4(55.6) 175.3(36.2) 163.0(31.7)
Neural A* 76.0(16.2) 64.7(27.7) 345.6(61.6) 306.5(46.9) 68.2(9.7) 37.6(21.7) 752.3(435.7) 727.0(423.1) 74.0(15.1) 53.3(57.7) 194.2(41.9) 182.9(36.9)
iA*(Ours) 80.0(13.4) 69.7(23.4) 342.9(62.0) 307.3(48.4) 71.0(12.2) 39.5(28.2) 738.0(435.6) 711.8(423.6) 71.2(12.0) 46.2(60.7) 171.4(35.2) 159.5(30.3)
IV-B Metrics
  • Reduction ratio of node explorations (Exp) [yonetani2021nastar]: To measure the search efficiency of the examined method, we utilize the ratio of the reduced search area to the classical A* search as a metric, which is defined as
    $\textit{Exp}=100\times\frac{A^{*}-A}{A^{*}},$ (10)
    where $A$ represents the search area of the examined method and $A^{*}$ is the search area of classical A* search.
  • Reduction ratio of time (Rt): To measure the time efficiency of the examined method, we utilize the ratio of saved time to the operation time of standard A* search as a metric, which is defined as
    $\textit{Rt}=100\times\frac{t^{*}-t}{t^{*}},$ (11)
    where $t^{*}$ is the operating time of classical A* solution, and $t$ is the operating time of the examined method.
    $\textit{AL}=\sqrt{\mathcal{L}_{a}(\mu)}+\mathcal{L}_{l}(\mu),$ (12)
    where $\mathcal{L}_{a}(\mu)$ and $\mathcal{L}_{l}(\mu)$ are the search area and path length, given path $\mu$. For a consistent unit, we use the square root of $\mathcal{L}_{a}(\mu)$ plus $\mathcal{L}_{l}(\mu)$ as the final metric.
  • Path Length (PL): To measure the path optimality of the examined method, we directly use the average of the absolute path length $\mathcal{L}_{l}(\mu)$ as a metric.
IV-C Overall Performance
@@
Comparisons with Benchmarks: To evaluate the path planning performance of iA*, we conduct both qualitative and quantitative experiments. Fig. 4 visualizes the planning results of the selected methods. From the results, the proposed iA* can greatly reduce the search area and still generate paths similar to optimal paths generated by classical A*. Furthermore, Table I summarizes the quantitative experimental results across the MP, Maze, and Matterport3D datasets with three different map sizes $64\times 64$, $128\times 128$, and $256\times 256$. To ensure fairness, we conducted 10 randomized trials, totally including 80 instances, for each kind of map in all three datasets, sampling start and goal positions uniformly. Final results are the average scores of these trials with the mentioned metrics. Notably, the transformer-based FS + PPM and WA* + CF support only maps with a fixed shape, $64\times 64$, because their network designs include fully connected layers that necessitate a predetermined input shape. Adjustments and re-training procedures are required to meet varied input instance shapes. Thus, comparisons at larger map sizes mainly focus on JPS, WA*, Neural A*, and iA*. Since there is no available batched JPS algorithm running on GPU, we exclude the $Rt$ metric in JPS results for fair comparisons. In the path length comparison, the results of the JPS algorithm are excluded, as JPS is designed to guarantee path optimality when using an admissible heuristic.
From the quantitative experimental results on the MP dataset, it is evident that the JPS algorithm exhibits excellent performance in reducing the search area by approximately 98% while maintaining path optimality. However, in the Matterport3D and Maze datasets, JPS performs poorly since jumps are constrained by frequent walls or turns, resulting in their poor performance on Exp and AL. The FS + PPM algorithm faces the similar limitation: it performs well on the MP dataset but deteriorates in the other two datasets. Different from the JPS algorithm, FS + PPM is likely overfitting on the MP dataset due to its powerful embedded neural network. The rest of the selected methods, WA*, Neural A*, WA* + CF, and iA* demonstrate consistent performance among the diverse maps with different sizes.
@@
iA* has satisfactory performance among them. The proposed iA* gets the lowest AL scores in 6 comparisons on 9 diverse environments with different map sizes. iA* saves on average reduction of 65.7% in search area, 54.4% in search time while still keeping the suboptimal paths with only a 3.8% increase on path length. Compared with the baseline Neural A* , iA* reduces the average search area by 3.3% and saves the runtime by 11.5%. Meanwhile, it achieves a 6.3% improvement in the trade-off between path length and search area, alongside a 6.1% reduction in path length. These results demonstrate that iA* can effectively minimize the search area while still identifying suboptimal paths. Notably, its consistent performance across different map sizes validates the method’s strong generalization without overfitting, which is critical for robotic applications.
@@
A* Neural A* iA*
Indoor
Exp (Area) 0 (9826) 3.95% (9438) 43.28% (5573)
@@
PL 801.02 822.23 809.40
IV-D Autonomous Exploration Development Environments
Applying the proposed iA* in real world needs a complete environment map, a reliable localization algorithm, and a stable wheeled platform. Thus, to validate the practicality, we evaluate iA* and baseline methods in the widely used simulation Autonomous Exploration Development Environment  [cao2022Autonomous] that includes a mobile robot and several scenarios resembling real-world settings. We use the indoor and the tunnel scenarios, shown as Fig. 5 (a), as our test environments. Specifically, we convert their dense point cloud maps into two-dimensional binary maps, where each grid cell represents a $0.5\mathrm{m}\times 0.5\mathrm{m}$ area in the original map. A grid cell will be regarded as an obstacle if its maximum $z$-value exceeds a given threshold; otherwise, it will be regarded as a free cell. Ultimately, the Tunnel and Indoor scenes have a size of $654\times 505$ and $259\times 199$, respectively. Since these map sizes are different, Neural A* is the only applicable data-driven baseline method. As a result, we mainly compare the performance of A*, Neural A*, and iA*. It is worth noting that both the Neural A* and iA* use the MP dataset for training and never see these scenarios.
In tests, the goal position for the mobile robot is manually given. We use the selected methods to generate paths and utilize the integrated path-following method within the maps to provide control commands for navigation. As shown in Fig. 5, both Neural A* and iA* can generate obstacle-free paths, similar to the optimal path generated by A*, and allow the robot to finish the task successfully. As shown in Table II, iA* significantly improves the planning efficiency compared to the baseline methods. Specifically, in the Indoor scenario, iA* reduces 43.28% search area and saves 52.04% time; in the Tunnel scenario, it reduces 62.73% search area and saves 62.84% time. Besides, in these two scenarios, iA* gets lower AL values, compared with Neural A*, presenting a better trade-off between search area and path length.
As the map size increases by 6.4 times (from 259×199 to 654×505), traditional A* exhibits a 4.6-fold growth in search area and runtime, while iA* scales more efficiently with only a 3.3-fold increase. This highlights iA*’s superior adaptability to larger environments, maintaining faster computation.
@@ IV-E Ablation Study
Figure 5: The trajectories of a mobile robot under given navigation tasks in the simulation environments, Tunnel (the first row) and Indoor (the second row). (a) represents the given instances, start (red circle), and goal (green circle). (b), (c) and (d) represent the trajectories of a mobile robot navigating from start to goal in the simulation environments with Classical A*, Neural A*, and iA*.
(a) Ablation study on imperative learning. (a) - (d) record the AL curves of four different configurations in validation process with and without imperative learning. The four configurations are formed by pairing two search engines (A* and WA*) and two network architectures (CNN and UNet) The solid lines denote the methods with IL-based training, while dashed lines are those without IL.
@@
We next evaluate the effectiveness of the IL strategy under different configurations. Four different configurations are considered by combining two search engines (A* and WA*) and two network architectures (CNN and UNet). Each configuration is trained both with and without the proposed IL training strategy. For non-IL training methods, we follow the Neural A* framework, employing the ground truth optimal paths as supervision. All other training settings, including the learning rate and optimizer, are kept consistent to ensure a fair comparison. To assess the trade-off between path optimality and search efficiency, we track the variation of the AL throughout the validation processes and show it in Fig. LABEL:fig:AUNET-LABEL:fig:WACNN, where solid and dashed lines represent the methods with and without IL-based training, respectively. As illustrated, the methods with IL consistently converge to significantly lower AL values than their non-IL counterparts. On average, our IL-based methods achieve approximately a 9.79% reduction in AL, demonstrating the effectiveness of IL and balancing between path optimality and search efficiency.
@@
To further analyze the individual contributions of search engines (A* and WA*) and network architectures (CNN and UNet) under IL strategy, we isolate and compare the results (solid lines in Fig. 9(a)) of the IL-trained configurations. The results show that models using the A* search engine (red and blue curves) outperform those employing WA* (green and yellow curves) in terms of lower AL scores. Moreover, the UNet-based models present superior performance compared to CNN-based models when paired with the same search engine. Consequently, based on these empirical observations, we adopt the configuration comprising the A* search algorithm and the UNet architecture for our final framework.
V Conclusion and Future Work